Samuel and Jackson are brothers. Samuel is three years older than twice the age of
Jackson. If the sum of their ages is 18, how old are the brothers?

Answer by ikleyn(52781)

.

    x + (2x+3) = 18

    x + 2x = 18 - 3

    3x     = 15

     x     = 15/3

ANSWER.   The brothers are  5 and 18-5 = 13 years old.

Solved.